# PATCH: a deep learning method to assess heterogeneity of artistic practice in historical paintings

### v2

@@
Abstract
@@
The history of art has seen significant shifts in the manner in which artworks are created, driven by changes in technology and the art market. Thus, understanding creative processes is a central question in technical art history. In the Renaissance and Early Modern period, paintings were largely produced by master painters who directed workshops of apprentices and others who often contributed to the work produced. The masters varied significantly in both their artistic and managerial styles, meaning different combinations of artists and implements might be seen both between masters and within workshops or even individual canvases. Information about the way different workshops were managed and the processes by which artworks were created remains elusive, however. Machine learning methods have potential to unearth new information about artists’ creative processes by extending the analysis of brushwork to a microscopic scale. Analysis of paintings created in a workshop setting, however, presents a challenge in that documentation of the artists and materials involved in individual works is sparse, meaning external examples are not available to train networks to recognize their contributions. Here we present a novel machine learning approach we call pairwise assignment training for classifying heterogeneity (PATCH) that is capable of identifying individual artistic practice regimes with no external training data, or “ground truth.” The method achieves unsupervised results by supervised means, and outperforms both simple statistical procedures and unsupervised machine learning methods. We apply this method to two historical paintings by the Spanish Renaissance master, El Greco: The Baptism of Christ and Christ on the Cross with Landscape, and our findings regarding the former potentially challenge previous work that has assigned the painting to workshop members. Further, the results of our analyses create a measure of heterogeneity of artistic practice that can be used to characterize artworks across time and space.
Introduction
@@
There are many ways to create a painting. The process can be collaborative or solitary, and there is no end to the styles, tools, materials and techniques that can be implemented. The evolution of what we call artistic practice—the technical, stylistic, material, and even physiological aspects of an artist’s creative process—over time and its variation between artists remains a key element of technical art history scholarship. For example, in the last two centuries, Western painters, aided (and in many ways, enabled) by concomitant changes in the tools of the trade such as the invention of the paint tube1, have often chosen more solitary practice in the studio or even outdoors. From the late Middle Ages through the Early Modern period, on the other hand, painting was typically done in a bustling workshop2. A master painter would establish a workshop and assemble a group of apprentices, journeymen and specialist painters and even collaborate with associates to help fulfill commissions and create smaller works to support the enterprise between larger contracts.
@@
Workshop practice adds a second layer of complexity to artistic practice. Master painters varied considerably in their managerial styles. Some performed much of the work themselves, while others dictated the composition but left many elements to the hands of workshop members (e.g., van Dyck3). Still others involved apprentices in the conceptualization of a piece4. Often, there was a hierarchy, with young apprentices typically assigned more mundane tasks such as the mixing of paints and the construction of canvases and more senior workshop members directed to paint portions of larger works in the master’s style5.
@@
Because workshops varied so widely in their organization and management yet produced works with cohesive styles, deeper understanding of variation in artistic practice can shed light on workshop practice and vice versa. Research into these key aspects of Renaissance and Early Modern art production is hindered by a lack of surviving textual sources regarding the structure, size, and membership of individual workshops as well as their day-to-day operation and the variation in artistic practices5. While scholars have been able to reconstruct individual artists’ workshops through a variety of methods, our collective understanding of the inner workings of workshops is still fragmentary6. Artistic practice varies widely, encompassing materials, methods, styles, and individual contributions. Understanding who interacted with individual paintings and how they did so—discerning the artists at work in a painting, the interplay of their individual “hands,” and the interaction of the artist’s hand and eye with paint and brush7—is an important open question in art history8. Answering these questions requires new methods of analysis.
@@
Machine learning has recently emerged as a promising complement to traditional art historical analyses in applications including preservation, conservation and the detection of forgeries9–12 as well as in fighting illegal trafficking in antiquities13. Much work involving style is focused on broad classification of works14,15, such as the use of a neural network to arrange works of the Western canon in the correct chronological order16. Recently published work by some of the authors of the present study demonstrated the efficacy of ML for attribution of paintings by analyzing high-resolution topographic images of the surface texture of oil paintings by known authors (student painters) in a controlled experimental setting. The topographic images, created using optical profilometry, captured variations in the height of the paint with microscopic resolution. While existing methods of study and conservation can employ microscopic methods to small areas obtained using a variety of imaging techniques, optical profilometry is unique in that it creates a topographic image at microscopic resolution over the entire painting. Though time consuming, the method provides a wealth of data on the surface topography and even the topography of pigmented layers beneath a transparent varnish layer. In a supervised machine learning analysis, a convolutional neural network (CNN) was trained to sort microscopic surface textures among artists, assigning 1 cm2 patches of paintings to the correct authors with 95% accuracy17. Thus, the method has shown that microscopic analysis can complement traditional art historical analyses of macroscopic features.
@@
But can deep learning on topographic images be used for such analyses when there is no ground truth to test against? When the number of artistic practice regimes in a given piece is unknown, known examples of the included artists’ work may not exist, and knowledge of the painting practices employed is incomplete—as in a painting produced by a Renaissance or Early Modern artist’s workshop? Here, we demonstrate a novel technique, to our knowledge, for ML-based attribution of paintings by unknown authors to the correct artist(s). We then combine this method with network analysis to identify artistic practice regimes (combinations of artists and materials) and create a measure of the heterogeneity of artistic practice (HAP) within a given painting or set of paintings. We apply this method to two important historical paintings by El Greco, one which is considered to be entirely by the master himself, and one previously thought to feature the work of members of his workshop.
ML methods can be divided into two classes based on the way the network is trained and the output of the network: supervised and unsupervised. In unsupervised learning, the machine analyzes individual objects (e.g., images or patches of an image) and creates classes based on the statistical properties of those objects. In supervised learning, the classes are already known, and the network learns by attempting to correctly classify objects in a training set, adjusting the weights of the interconnections between its nodes as it succeeds (or fails) in doing so. A successfully trained network will be able to assign novel objects to the correct classes. Supervised learning is thus particularly useful for image recognition tasks. The method employed in the previous publication described above17 was a standard application of supervised learning for image recognition to topographic images. The network was trained on known topographic images (in this case, two paintings by each artist) and then tested on novel objects (a third painting by each artist). The trained network was able to distinguish between four different artists, correctly assigning 1 cm2 patches of each painting to its respective artist with 95% accuracy.
@@
The Pairwise Assignment Training for Classifying Heterogeneity (PATCH) method described here uses supervised learning toward an unsupervised end. Rather than train a network to recognize individual classes, we test whether the network is capable of learning to distinguish between two objects and then use that information to build post-hoc classes. The PATCH method relies on the inability of the network to correctly sort patches of paintings or regions of a painting painted by the same artist. Given patches of two paintings by different painters—a different-artist pair—the network will learn to assign those patches to the correct painting with a high degree of accuracy, as demonstrated in our previous publication. But if the network is given patches of two paintings by the same artist—a same-artist pair—it will not be able to learn to assign those patches to the correct painting; it will perform no better than a coin flip. We then create a network of same-artist pairs and use network analysis to construct classes corresponding to artistic practice regimes. In this way, we obviate the need for separate ground truth information for individual artists and artistic practice regimes, enabling successful applications in contexts such as workshop paintings where the ground truth is not known with certainty.
Study 1: Method development
@@
The first phase was developed on a set of paintings by known artists. A group of nine artists each individually painted three paintings, all using the same tools, materials, and subject (a photograph of lilies). The materials used for the paintings were described previously17. While the artists were instructed to realistically depict the subject, choices of style and technique (including the use of brushes and alteration of paint viscosity) were left up to the artists. One artist was only able to produce two paintings, and one painting was removed from the sample due to the optical profilometry data being corrupted, for a total of 25 paintings. Topographical information for each painting was recorded using high resolution spectral confocal optical profilometry with a spatial resolution of 50 microns, and a height repeatability of 200 nm. The resulting data were processed to remove any large scale warping of the canvas by subtracting a mean-filtered version of the height map (with filter radius of 0.5 cm)17.
@@
The network architecture chosen was VGG-16, implemented in TensorFlow and pretrained on the ImageNet dataset. Given the computational complexity of the training—the combinatorics of many region-to-region comparisons, repeated over many folds, described in detail below—the relatively light-weight but accurate VGG-16 architecture was ideal for the PATCH approach. Output from VGG-16 was pooled and flattened and passed through two dense layers with dropout rates of 0.25. The learning rate was set to 0.0001 and the batch size was set to 32 patches (see the Supplementary Materials for more information on hyperparameter optimization). In order to measure the ability of the network to learn to distinguish between two paintings or regions of a painting, we ran 26 training folds. Each fold comprised the following steps: 1 cm2 patches from each painting/region were randomly selected, with replacement. Sampling with replacement creates a bootstrapping effect that drives down the success rate of same-artist pairs and drives up the success rate for different-artist pairs. In each fold, the network was trained for 25 epochs with validation set size set to 30% of patches, and the maximum validation accuracy (percentage of validation patches correctly identified) was recorded. At the end of 26 folds, the distribution of the 26 maximum validation accuracies was compared to the distribution expected if the network were assigning patches at random.
@@
If the network has failed to learn to assign objects to their correct classes, it should perform this task no better than if it were assigning them randomly. The prospect of randomly choosing from two options invokes the coin flip and, by extension, the binomial distribution. However, we are measuring the maximum validation accuracy from 25 training epochs in each of 26 folds. Thus, the distribution expected if the network fails to learn to distinguish between painters would be the random assignment distribution (RAD) , the probability that you will see a maximum of m heads (i.e., correct assignments) over k repetitions (k epochs) of an experiment where you flip a fair coin n times (the number of patches in the test set). The Supplementary Materials contains an analytical derivation of this distribution.
@@
Two decision criteria were selected to determine whether the distribution of maximum validation accuracies indicated that the network had failed to learn to distinguish between the input paintings: 1) if the observed mean had a z-score less than 2 relative to the RAD mean and 2) if the largest of the observed maximum validation accuracies (the right edge of the observed distribution) among the 26 folds was less than an empirically determined threshold based on finite sampling from RAD, plus 10% to account for the effects of bootstrapping (selection with replacement, see Supplementary Information for more details). Using these decision criteria, the accuracy in determining whether the artist was the same or different was exceptional. We identified same-artist pairs (n = 23) and different-artist pairs (n = 277) with 96.0% overall accuracy (87% for same and 99.3% for different). This translates to F1 scores of 0.991 for different-artist pairs and 0.889 for same-artist pairs, an average of 0.940 (Table 1). The pairwise assignment training phase outperformed both an unsupervised method and a statistical method based on surface roughness (see details in SI Appendix C) in precision, recall and total F1 for both same-artist and different-artist pairs.
| Precision | Recall | F1 | |
| Same-artist pairs | 0.909 | 0.87 | 0.889 |
| Different-artist pairs | 0.989 | 0.993 | 0.991 |
| Average | 0.949 | 0.931 | 0.940 |
@@
While pairwise assignment training performs exceptionally well in identifying artists, situations where ground truth about the artistic practices employed in a particular painting is not available require some means to create the post hoc classes to which regions of a painting will be assigned. To this end, we use network analysis. We construct a network where same-artist pairs (or, for historical paintings, same-practice pairs) of paintings or regions are connected by an edge (lines connecting regions in Figure 3) and different-artist/practice pairs are unconnected.
@@
With potentially spurious edges removed, and the remaining edges weighted equally, we then employ a community finding algorithm to identify groups of regions that are the most similar. If a set of regions has a large number of internal links within the set, and fewer links to outside regions, it forms a “community” within the network18–20, suggesting that the regions in a given community were likely painted under different circumstances (artists, materials, etc.) than those in other communities. To establish the degree of difference between communities, we can characterize the community structure of the network via measures such as modularity ()—the fraction of edges in the network that are internal to communities minus the mean fraction in a network with the same communities but where the edges are completely randomized19. We implemented the Louvain community finding algorithm21 in Gephi22. With the resolution set to 1.0, the algorithm returns the partition of the network with the maximum modularity. The Louvain algorithm returned a correct partition of the experimental dataset with .
@@
El Greco, born Domenikos Theotokópoulos (1541-1614), is regarded as a pillar of the Renaissance in Spain and as an early progenitor of modernism23. El Greco became a master of icon painting in Crete before journeying through Venice (where he may have worked with Titian) and Rome, eventually settling in Toledo, Spain24,25. It was in Toledo that the master married Byzantine and Venetian motifs26 and developed the peculiar style for which he is most widely recognized, with its distorted human figures, “expressive hands”27, and ”exploitation” of ”pure colors to their limits”28. Two examples of that style are The Baptism of Christ (1624, Hospital Tavera, Toledo, Esp., henceforth, The Baptism), shown in Figure 2A, and Christ on the Cross with Landscape (ca. 1600-1610, Cleveland Museum of Art, Cleveland, OH, henceforth Christ on the Cross), shown in Figure 2B.
@@
These works were chosen for this analysis because of the difference in how art historians have characterized their authorship. The Baptism has long been thought to feature the work of El Greco and at least one other artist. Primary historical evidence indicates that El Greco began the painting under a contract with the Hospital Tavera but retained it at his death in 1614. It was delivered to the Hospital nearly a decade later29, during which period art historians have proposed the painting was finished by workshop members, particularly the master’s son, Jorge Manuel24. Previous art historical studies have attempted to attribute regions of The Baptism to the master and others (illustrated in Figure S2) largely through connoisseurship—visual analysis of artists’ styles and artistic choices30. El Greco himself is proposed to have painted the entire top, with the possible exception of the robe of the angel on the right (which was not scanned for this study due to its poor condition)29,31, as well as the angel in green (bottom left), with the exception of the wings. Jorge Manuel is proposed to have been painted John the Baptist (bottom right) and the adjacent figure in red. Lopera31 and Wethey29 ascribe the image of Christ (bottom center) to Jorge Manuel and El Greco, respectively, suggesting that it could represent the work of both artists. Lopera believes that a third hand may have been involved in rendering the faces of the angels in the background on the bottom. The landscapes at the bottom were extended and “transformed into a river” by an unknown artist sometime between receipt of the painting and its installation in the epistle-side altarpiece ca. 1660, which was not an unusual practice at the time. Restoration of the landscape on the right at the very bottom was performed in 193629,31,32. In contrast, Christ on the Cross has been entirely attributed to El Greco himself. It is generally accepted that El Greco created many works in their entirety without the aid of his workshop, including Christ on the Cross. Francis33 states that most of El Greco’s approximately twenty renditions of the Crucifixion were “by his own hand,” and notes the “lightning dexterity of brushwork” in this particular example.
Study 2: Application to El Greco’s The Baptism
@@
Two large sections of The Baptism were scanned in situ using the Factum Foundation’s Lucida scanner. The Lucida scanner is a non-contact laser triangulation scanner with 100-micron lateral resolution and is capable of acquiring scans in situ over a large area by stitching together 48 cm square tiles. The scanned regions used in this study are indicated in Figure 3. The height data was processed in the same way as in study 1, including correction for possible canvas warping.
@@
Regions of the topographic image of The Baptism were selected by hand in order to minimize variation in subject matter (e.g., faces, robes) within each region and to constrain region size to between 180 and 540 cm2. The lower bound was chosen because the method had been developed using 180-patch paintings in study 1 and the upper bound was selected to ensure adequate representation of patches when regions of significantly different size were compared. We avoided selecting areas with significant cracking or damage, including the robes of the large angels to the left and right of God, and the red swatch held by the “green angel” in the bottom half of the painting. Each fold proceeded as described in study 1, above. Where regions were of different sizes, the sample size was set equal to the number of patches in the smaller region.
@@
Pairwise comparisons of all 55 regions (a total of 1485 pairwise tests) of The Baptism yielded a network with 356 edges, which was reduced to 314 edges after pruning. The Louvain community finding algorithm21 with resolution set to 1.0, returned a partition with that featured four communities (Figure 3a). A value of is often considered evidence of structure in real-world networks19.
@@
Study 3: Application to El Greco’s Christ on the Cross
As in study 2, regions of the topographic image of Christ on the Cross were selected by hand in order to minimize variation in subject matter and to constrain region size. Areas with apparent cracking or damage, such as the background and trees on the lower left and the cross beam to the immediate left of Christ’s face, were not included.
Pairwise comparison of all 24 regions (276 pairwise tests) yielded a network with 123 edges, which was reduced to 108 edges after pruning. The Louvain community finding algorithm21 at resolution 1.0 generated a partition with that featured two communities. Two is the smallest possible number of communities for a non-fully connected graph, as compares inter- and intra-community connections. Communities are mapped onto the painting in Figure 4. Constituent regions of both communities are not randomly distributed, with one community overlapping much of Christ’s torso and face, while the other community overlaps Christ’s legs, the lower portion of the cross and regions of the background.
Analysis of PATCH-generated communities
Further analysis of the community structure in the PATCH generated networks can provide additional insight into the community cohesiveness beyond the modularity index. We analyzed the average community degree, defined as the count of all edges between nodes within two communities. This is shown for all possible community pairs in Figure 5A and C, and within a single community (internal degree) or between nodes within and external to a given community (external degree) in Figure 5B and D. Each degree is normalized by the theoretical maximum degree based on the number of edges in the classification. The degree between community pairs and external degree can be taken as a measure of the PATCH inter-community similarity between specific communities and between a community and all others, respectively. Likewise, the internal degree can be taken as a measure of intra-community cohesiveness.
In The Baptism, all of the connections between the first three communities are significantly stronger than those involving the fourth community, with the average external degree over four times larger. This indicates a significant isolation of the fourth community (colored orange in Figure 3, and represented by orange bars in Figure 5B), indicating a relatively strong disparity in the topographical data. The strongest connection is between communities 1 (yellow) and 3 (green) and the weakest is between 3 and 4, which has no connecting edges. Overall the 3rd community has a slightly greater connection with the remainder of the communities, indicating greater similarity. Interestingly, the internal degree is relatively similar for all communities, indicating that the differences within the observed topographical signal are evenly distributed across communities. A significant difference in the internal degree could, for instance, indicate a community with above or below average range of stylistic practice.
Although the external degree analysis is largely redundant in the case of Christ on the Cross, since only two communities are proposed, it should be noted that the external degree between the two communities is significantly larger than for any other two communities in The Baptism except the connection between communities 1 and 3. This indicates the two proposed communities are on average more closely connected than those proposed in The Baptism, which is reflected in the low modularity value. Additionally, the internal degree in each community is similar, which can indicate a similar intra-community stylistic distribution.

### v3

@@
PATCH: a deep learning method to assess heterogeneity of artistic practice in historical paintings
@@
In the Renaissance and Early Modern period, paintings were largely produced by master painters who directed workshops of apprentices and others who often contributed to the piece. Discerning who created these masterworks and how they did so is a central question in technical art history1 and a nontrivial problem that machine learning can help solve by extending analysis to a microscopic scale2. Analysis of workshop paintings presents a challenge, however, because information about the members of workshops and the processes by which artworks were created remains elusive3; 4. Thus, external examples are not available to train networks to recognize. Here we present a novel machine learning approach we call pairwise assignment training for classifying heterogeneity (PATCH) that is capable of identifying individual artistic practice regimes with no external training data. We apply this method to two historical paintings by the Spanish Renaissance master, El Greco, and our findings regarding one of the works potentially challenge previous studies that assert that a considerable portion of the painting was completed by workshop members after El Greco’s death5. PATCH outperforms statistical and unsupervised machine learning methods in this complex pairwise comparison problem lacking “ground truth” data, making it potentially useful across similar cases in the social and natural sciences, including image segmentation in remote sensing6, urban development and design7, and anomaly detection manufacturing contexts8; 9, among others.
@@
There are many ways to create a painting. The process can be collaborative or solitary, and there is no end to the styles, tools, materials and techniques that can be implemented. The evolution of what we call artistic practice—the technical, stylistic, material, and even physiological aspects of an artist’s creative process—over time and its variation between artists remains a key element of technical art history scholarship. From the late Middle Ages through the Early Modern period, on the other hand, painting was typically done in a bustling workshop10. A master painter would establish a workshop and assemble a group of apprentices, journeymen and specialist painters and even collaborate with associates to help fulfill commissions and create smaller works to support the enterprise between larger contracts.
@@
Workshop practice adds a second layer of complexity to artistic practice. Master painters varied considerably in their managerial styles. Some performed much of the work themselves, while others dictated the composition but left many elements to the hands of workshop members (e.g., van Dyck11). Still others involved apprentices in the conceptualization of a piece12. Often, there was a hierarchy, with young apprentices typically assigned more mundane tasks such as the mixing of paints and the construction of canvases and more senior workshop members directed to paint portions of larger works in the master’s style3.
@@
Because workshops varied so widely in their organization and management yet produced works with cohesive styles, deeper understanding of variation in artistic practice can shed light on workshop practice and vice versa. Research into these key aspects of Renaissance and Early Modern art production is hindered by a lack of surviving textual sources regarding the structure, size, and membership of individual workshops as well as their day-to-day operation and the variation in artistic practices3. While scholars have been able to reconstruct individual artists’ workshops through a variety of methods, our collective understanding of the inner workings of workshops is still fragmentary4. Artistic practice varies widely, encompassing materials, methods, styles, and individual contributions. Understanding who interacted with individual paintings and how they did so—discerning the artists at work in a painting, the interplay of their individual “hands,” and the interaction of the artist’s hand and eye with paint and brush13—is an important open question in art history1.
@@
Machine learning has recently emerged as a promising complement to traditional art historical analyses in applications including preservation, conservation, and the detection of forgeries14; 15; 16; 17, as well as in fighting illegal trafficking in antiquities18. Much work involving style is focused on broad classification of works19; 20, such as the use of a neural network to arrange works of the Western canon in the correct chronological order21. Recently published work by some of the authors of the present study demonstrated the efficacy of ML for attribution of paintings by analyzing high-resolution topographic images of the surface texture of oil paintings by known authors (student painters) in a controlled experimental setting. In a supervised machine learning analysis, a convolutional neural network (CNN) was trained to sort microscopic surface textures among artists, assigning 1 cm2 patches of paintings to the correct authors with 95% accuracy2. Thus, the method has shown that microscopic analysis can complement traditional art historical analyses of macroscopic features.
Unfortunately, supervised deep learning on topographic images cannot be used when there is no ground truth for the network to learn on. When the object of analysis is a painting produced by a Renaissance or Early Modern artist’s workshop, the number of artistic practice regimes is generally unknown, known examples of the included artists’ work may not exist, and knowledge of the painting practices employed is incomplete. Unsupervised learning, wherein the machine analyzes individual objects (e.g., images or patches of an image) and creates classes based on the statistical properties of those objects, would appear more appropriate. However, a naive unsupervised analysis still depends on hyperparameters that can be tuned to yield larger or smaller numbers of classes. Whether performed by the trained art historian or by a neural network, the problem of identifying regions of shared artistic practice is, in a word, nontrivial.
@@
Here, we demonstrate a novel technique, to our knowledge, for ML-based image attribution, and use it to assign paintings to their respective artist(s). We then combine this method with network analysis to identify artistic practice regimes (combinations of artists and materials) and create a measure of the heterogeneity of artistic practice (HAP) within a given painting or set of paintings. We apply this method to two important historical paintings by El Greco, one considered to be entirely by the master himself, and one previously thought to feature the work of members of his workshop.
@@
The method described here, Pairwise Assignment Training for Classifying Heterogeneity (PATCH), uses supervised learning toward an unsupervised end. Rather than train a network to recognize individual classes, we test whether the network is capable of learning to distinguish between two objects and then use that information to build post-hoc classes. The PATCH method relies on the inability of the network to correctly sort patches of paintings or regions of a painting painted by the same artist. Given patches of two paintings by different painters—a different-artist pair—the network will learn to assign those patches to the correct painting with a high degree of accuracy, as demonstrated in our previous publication. But if the network is given patches of two paintings by the same artist—a same-artist pair—it will not be able to learn to assign those patches to the correct painting; it will perform no better than a coin flip. We then create a network of same-artist pairs and use network analysis to construct classes corresponding to artistic practice regimes. In this way, we obviate the need for separate ground truth information for individual artists and artistic practice regimes, enabling successful applications in contexts such as workshop paintings where the ground truth is not known with certainty.
Development and validation of the PATCH algorithm
@@
The first phase was developed on a set of paintings by nine known student artists. Each artist individually painted three paintings, all using the same tools, materials, and subject (a photograph of lilies), as described previously2. While the artists were instructed to realistically depict the subject, choices of style and technique were left up to the artists. One artist was only able to produce two paintings, and one painting was removed from the sample due to the optical profilometry data being corrupted, for a total of 25 paintings. Topographical information for each painting was recorded using high resolution spectral confocal optical profilometry with a spatial resolution of 50 microns, and a height repeatability of 200 nm. The resulting data were processed to remove any large-scale warping of the canvas by subtracting a mean-filtered version of the height map (with filter radius of 0.5 cm)2. Because in this dataset the paintings were small (12 cm 15 cm), each “region” chosen for the pairwise training consists of an entire painting. The region is then subdivided into 1 cm2 square patches, which corresponds to the size where the earlier supervised machine learning analysis 2 was able to achieve maximum classification accuracy. This patch size is big enough to contain significant information about brushstrokes, while yielding sufficiently large datasets for training (in this case 180 patches per painting).
@@
The network architecture chosen was VGG-16, implemented in TensorFlow and pretrained on the ImageNet dataset. Given the computational complexity of the training—the combinatorics of many region-to-region comparisons, repeated over many folds, described in detail below—the relatively light-weight but accurate VGG-16 architecture was ideal for the PATCH approach (see the Supplementary Materials for more information on hyperparameter settings and optimization). In order to measure the ability of the network to learn to distinguish between two paintings or regions of a painting, we ran 26 training folds. Each fold comprised the following steps: 1 cm2 patches from each painting/region were randomly selected, with replacement. Sampling with replacement creates a bootstrapping effect that drives down the success rate of same-artist pairs and drives up the success rate for different-artist pairs. To eliminate the influence of directional elements, the corners of each patch were trimmed to create an octagon, and the patch was randomly rotated to one of eight possible orientations. The network was then trained for 25 epochs with validation set size set to 30% of patches, and the maximum validation accuracy (percentage of validation patches correctly identified) was recorded. The distribution of the 26 maximum validation accuracies was then compared to the distribution expected if the network were assigning patches at random.
@@
If the network has failed to learn to assign objects to their correct classes, it should perform this task no better than if it were assigning them randomly. We are measuring the maximum validation accuracy from 25 training epochs in each of 26 folds. Thus, the distribution expected if the network fails to learn to distinguish between painters would be the distribution , the probability that you will see a maximum of m heads (i.e., correct assignments) over k repetitions (k epochs) of an experiment where you flip a fair coin n times (the number of patches in the test set), which we call the random assignment distribution (RAD) (see the Supplementary Materials for an analytical derivation of this distribution).
@@
Two decision criteria were selected to determine whether the distribution of maximum validation accuracies indicated that the network had failed to learn to distinguish between the input paintings: 1) if the observed mean had a z-score less than 2 relative to the RAD mean and 2) if the largest of the observed maximum validation accuracies (the right edge of the observed distribution) among the 26 folds was less than an empirically determined threshold based on finite sampling from RAD, plus 10% to account for the effects of bootstrapping (selection with replacement, see Supplementary Information for more details). Using these decision criteria, the classification performance in determining whether the artist was the same or different was exceptional. We identified same-artist pairs (n = 23) and different-artist pairs (n = 277) with F1 scores of .889 and .991 respectively (see Table S2). The PATCH algorithm substantially outperformed a statistical method based on surface roughness, and beat a variety of unsupervised and supervised clustering approaches (see details in SI Appendix C) in F1 for both same-artist and different-artist pairs.
@@
While pairwise assignment training performs exceptionally well in identifying artists with control for materials, situations where ground truth about the artistic practices employed is not available require some means to create the post hoc classes to which regions of a painting will be assigned. To this end, we use network analysis. We construct a network where same-artist pairs (or, for historical paintings, same-practice pairs) of paintings or regions are connected by an edge (lines connecting regions in Figure 3) and different-artist/practice pairs are unconnected.
@@
With potentially spurious edges removed and the remaining edges weighted equally, we then employ a community finding algorithm to identify groups of regions that are the most similar. If a set of regions has a large number of internal links within the set and fewer links to outside regions, it forms a “community” within the network22; 23; 24, suggesting that the regions in a given community were likely painted under different circumstances (artists, materials, etc.) than those in other communities. To establish the degree of difference between communities, we can characterize the community structure of the network via measures such as modularity ()—the fraction of edges in the network that are internal to communities minus the mean fraction in a network with the same communities but where the edges are completely randomized23. We implemented the Louvain community finding algorithm25 in Gephi26. With the resolution set to 1.0, the algorithm returns the partition of the network with the maximum modularity. The Louvain algorithm returned a correct partition of the experimental dataset (nine disjoint communities for nine different artists) with (see Figure S4).
@@
El Greco, born Domenikos Theotokópoulos (1541-1614), is regarded as a pillar of the Renaissance in Spain and as an early progenitor of modernism27. El Greco became a master of icon painting in Crete before journeying through Venice and Rome, eventually settling in Toledo, Spain5; 28. It was there that the master married Byzantine and Venetian motifs29 and developed the peculiar style for which he is most widely recognized, with its distorted human figures, “expressive hands”30, and ”exploitation” of ”pure colors to their limits”31. Two examples of that style are The Baptism of Christ (1624, Hospital Tavera, Toledo, Esp., henceforth, The Baptism), shown in Figure 2A, and Christ on the Cross with Landscape (ca. 1600-1610, Cleveland Museum of Art, Cleveland, OH, henceforth Christ on the Cross), shown in Figure 2B.
@@
These works were chosen for this analysis because of the difference in how art historians have characterized their authorship. The Baptism has long been thought to feature the work of El Greco and at least one other artist. Primary historical evidence indicates that El Greco began the painting under a contract with the Hospital Tavera but retained it at his death in 1614. It was delivered to the Hospital nearly a decade later32, during which period art historians have proposed the painting was finished by workshop members, particularly the master’s son, Jorge Manuel5. Previous art historical studies have attempted to attribute regions of The Baptism to the master and others (illustrated in Figure S3) largely through connoisseurship—visual analysis of artists’ styles and artistic choices33. El Greco himself is proposed to have painted the entire top, with the possible exception of the robe of the angel on the right (which was not scanned for this study due to its poor condition)32; 34, as well as the angel in green (bottom left), with the exception of the wings. Jorge Manuel is proposed to have painted John the Baptist (bottom right) and the adjacent figure in red. Lopera34 and Wethey32 ascribe the image of Christ (bottom center) to Jorge Manuel and El Greco, respectively, suggesting that it could represent the work of both artists. Lopera believes that a third hand may have been involved in rendering the faces of the angels in the background on the bottom. The landscapes at the bottom were extended and “transformed into a river” by an unknown artist sometime between receipt of the painting and its installation in the epistle-side altarpiece ca. 1660, which was not an unusual practice at the time. Restoration of the landscape on the right at the very bottom was performed in 193632; 34; 35. In contrast, Christ on the Cross has been entirely attributed to El Greco himself. It is generally accepted that El Greco created many works in their entirety without the aid of his workshop, including Christ on the Cross. Francis36 states that most of El Greco’s approximately twenty renditions of the Crucifixion were “by his own hand,” and notes the “lightning dexterity of brushwork” in this particular example.
Application to El Greco’s The Baptism
@@
Two large sections of The Baptism were scanned in situ using the Factum Foundation’s Lucida scanner. The Lucida scanner is a non-contact laser triangulation scanner with 100-micron lateral resolution and is capable of acquiring scans in situ over a large area by stitching together 48 cm square tiles. The scanned regions used in this study are indicated in Figure 3. The height data was processed in the same way as for the student paintings described above, including correction for possible canvas warping.
@@
Regions of the topographic image of The Baptism were selected by hand in order to minimize variation in subject matter (e.g., faces, robes) within each region and to constrain region size to between 180 and 540 cm2. The lower bound was chosen because the method had been developed and validated using 180-patch student paintings and the upper bound was selected to ensure adequate representation of patches when regions of significantly different size were compared. We avoided selecting areas with significant cracking or damage, including the robes of the large angels to the left and right of God, and the red swatch held by the “green angel” in the bottom half of the painting. When comparing regions of different sizes during the PATCH analysis, the sample size was set equal to the number of patches in the smaller region.
@@
Pairwise comparisons of all 55 regions (a total of 1485 pairwise tests) of The Baptism yielded a network with 356 edges, which was reduced to 314 edges after pruning. The Louvain community-finding algorithm25 with resolution set to 1.0, returned a partition with that featured four communities (Figure 3a). A value of is often considered evidence of structure in real-world networks23.
@@
Application to El Greco’s Christ on the Cross
As in the case of The Baptism, regions of the topographic image of Christ on the Cross were selected by hand in order to minimize variation in subject matter and to constrain region size. Areas with apparent cracking or damage, such as the background and trees on the lower left and the cross beam to the immediate left of Christ’s face, were not included.
Pairwise comparison of all 24 regions (276 pairwise tests) yielded a network with 123 edges, which was reduced to 108 edges after pruning. The Louvain community finding-algorithm25 at resolution 1.0 generated a partition with that featured two communities. Two is the smallest possible number of communities for a non-fully connected graph, as compares inter- and intra-community connections. Communities are mapped onto the painting in Figure 4. Constituent regions of both communities are not randomly distributed, with one community overlapping much of Christ’s torso and face, while the other community overlaps Christ’s legs, the lower portion of the cross and regions of the background.